# StreamDiT: Real-Time Streaming Text-to-Video Generation

### v1

@@ 4.2 Multistep Distillation
Sampling distillation is a key component to build real-time streaming video generation. Due to the modifications of frame partitioning and micro denosing steps made by StreamDiT, standard sampling distillation methods [35, 31, 14] cannot be applied. Thus, a customized sampling distillation for StreamDiT is needed.
As described in Sec. 3.2, StreamDiT partitions frames in a buffer to $N$ chunks, each of which has $s$ denoising steps before moving the buffer. In particular, we aim at reducing the micro-step $s$ in Eq. 7. Our StreamDiT model is trained with mixed partitioning schemes, to make it flexible to different scenarios. In distillation, we first choose a partitioning scheme. Specifically, we set $K=0$, $c=2$, $s=16$, $N=8$ for the teacher model. This indicates that the teacher model has $s*N=128$ denoising steps with CFG. The FM trajectory is split to $N$ segments. We then perform step distillation in each segment separately. In practice, we perform both step and guidance distillation at the same time, by distilling multiple CFG steps of the teacher into a single conditional forward pass of the student.
@@ 5.1 Implementation Details
We finetune a T2V model [33] with 4B parameters using the architecture introduced in Sec. 4.1. The latent size is $[16,64,64]$. With a [4×,8×,8×][4$\times$,8$\times$,8$\times$][ 4 × , 8 × , 8 × ] TAE, the base model generates 64 frames at 512p resolution.
@@
Subject Consistency Background Consistency Temporal Flickering Motion Smoothness Dynamic Degree Aesthetic Quality Quality Score
@@ 5.2 Evaluation
We compare our method with ReuseDiffuse [11] and FIFO-Diffusion [20] for streaming generation of long videos. To make a fair comparison, we implemented their methods on our base model, so they share the same visual quality. We select 50 prompts from the evaluation dataset of Movie Gen [33] that are suitable for long videos. We adopt VBench [19] quality metrics for quantitative evaluation.
@@
In addition to VBench metrics, we also conduct human evaluations following the guidance in [33]. A human evaluation compares two model results on the same prompts side-by-side. Annotators were asked to evaluate pairs of video samples to determine which one was superior or if they were of comparable quality along several axes. The evaluation criteria encompassed four aspects: overall quality, frame consistency, motion completeness, and motion naturalness. For these evaluations, we employ the same set of 50 prompts used in the VBench evaluation, generating 8-second videos at a resolution of 512p for each prompt. As shown in Fig. 5, our proposed method surpasses existing approaches across all evaluation metrics.
@@
Table 3: Ablation of mixed training. Chunk size 1 represents the Progressive AR Diffusion [42], and chunk size 16 represents the original T2V without streaming. A mixed training with all chunk sizes show the best quality score.
@@ 5.3 Ablation Study
We conduct an ablation study to analyze the mixed training discussed in Sec. 3.2. Specifically, we train models with different mixing schemes as shown in Tab. 3. Chunk size 1 represents the Progressive AR Diffusion [42] implemented on our model. Recall that our base model has 16 latent frames, so chunk size 16 is the basic T2V. After the models are trained, we generated videos using chunk size $c=1$, since this is the case covered by all models. Among them, we see that the mixture of all chunk sizes achieves the best quality score, although the inference is biased to chunk size 1 as being set in inference. Please note that we use a different inference scheme in Tab. 2, so the numbers of our model are different. The first one with chunk size 1 and no mixing achieves the second-best quality score, indicating an overfitting to this case.
@@ 5.4 Applications
Real-Time Streaming: Our distilled model has a chunk size 2 and 8 sampling steps without CFG. We benchmark its performance on one H100 GPU. It takes 482 ms for one denoising step to generate 2 latent frames and 8 video frames after a 4×4$\times$4 × TAE. The latency of text encodering and TAE decoding are negligible. Thus, our distilled model can reach real-time performance at 16 FPS. Fig. 7 shows frames from a one-minute long video generated in real-time. With real-time streaming generation, our model can potentially work as a game engine.
@@
Video-to-Video Streaming: Our model can also be applied to real-time video-to-video tasks. The video-to-video process is based on the noise addition and denoising strategy proposed in SDEdit [30]. First, we add noise to the original input video and then denoise the noisy video with new prompt guidance. We achieve significant content editing while maintaining good temporal consistency. The visual results shown in Fig. 10 highlight the model’s powerful video editing capabilities. In this example, a pig in the input video is modified to a cat in the output video, while the background is preserved.
@@ 7 Acknowledgment
We would like to express our sincere gratitude to Yen-Cheng Liu, Markos Georgopoulos, Felix Juefei-Xu, Luxin Zhang, and Tao Xu for their valuable discussions and insightful contributions, which greatly enriched our work.
@@ 10 More Results
To evaluate the capabilities in generating higher-quality videos, we further fine-tuned the 30B model from Movie Gen [33]. While the 30B model is not available for real-time generation at the present, it effectively supports the creation of exceptionally high-quality videos, as illustrated in Fig. 15. Our model consistently produces extended videos characterized by impressive visual quality, content coherence, and diverse scenes aligned accurately with provided text prompts.
@@ 11 Limitations
As described in Sec. 8, the effective context length of StreamDiT fundamentally depends on the base T2V model, and thus lacks long-term memory. When content falls outside the short-term memory window (i.e. context length) of StreamDiT, the associated information is likely to be lost. This can lead to issues such as identity mismatches in a person’s face, or background inconsistencies when the camera makes a full rotation. However, since StreamDiT is orthogonal to additional long-term memory mechanisms, it is possible to address this issue in the future by combining it with long-term memory architectures such as State Space Models like Mamba [10].

### v3

@@ 4.2 Multistep Distillation
Sampling distillation is a key component to build real-time streaming video generation. Due to the modifications of frame partitioning and micro denoising steps made by StreamDiT, standard sampling distillation methods [salimans2021progressive, guideddistillation, heek2024multistepconsistencymodels] cannot be applied. Thus, a customized sampling distillation for StreamDiT is needed.
As described in Sec. 3.2, StreamDiT partitions frames in a buffer to $N$ chunks, each of which has $s$ denoising steps before moving the buffer. In particular, we aim at reducing the micro-step $s$ in Eq. 7. Our StreamDiT model is trained with mixed partitioning schemes, to make it flexible to different scenarios. In distillation, we first choose a partitioning scheme. Specifically, we set $K=0$, $c=2$, $s=16$, $N=8$ for the teacher model. This indicates that the teacher model has sN=128s*N=128 denoising steps with CFG. The FM trajectory is split to $N$ segments. We then perform step distillation in each segment separately. In practice, we perform both step and guidance distillation at the same time, by distilling multiple CFG steps of the teacher into a single conditional forward pass of the student.
@@ 5.1 Implementation Details
We finetune a T2V model [moviegen] with 4B parameters using the architecture introduced in Sec. 4.1. The latent size is $[16,64,64]$. With a $4\times,8\times,8\times]$ TAE, the base model generates 64 frames at 512p resolution.
@@
Subject
Consistency
Background
Consistency
Temporal
Flickering
Motion
Smoothness
Dynamic
Degree
Aesthetic
Quality
Quality
Score
@@ 5.2 Evaluation
We compare our method with ReuseDiffuse [gu2023reusediffuse] and FIFO-Diffusion [Kim2024FIFO] for streaming generation of long videos. To make a fair comparison, we implemented their methods on our base model, so they share the same visual quality. We select 50 prompts from the evaluation dataset of Movie Gen [moviegen] that are suitable for long videos. We adopt VBench [huang2023vbench] quality metrics for quantitative evaluation.
@@
In addition to VBench metrics, we also conduct human evaluations following the guidance in [moviegen]. A human evaluation compares two model results on the same prompts side-by-side. Annotators were asked to evaluate pairs of video samples to determine which one was superior or if they were of comparable quality along several axes. The evaluation criteria encompassed four aspects: overall quality, frame consistency, motion completeness, and motion naturalness. For these evaluations, we employ the same set of 50 prompts used in the VBench evaluation, generating 8-second videos at a resolution of 512p for each prompt. As shown in Fig. 5, our proposed method surpasses existing approaches across all evaluation metrics.
@@
Table 3: Ablation of mixed training. Chunk size 1 represents the Progressive AR Diffusion [xie2024progressiveautoregressivevideodiffusion], and chunk size 16 represents the original T2V without streaming. A mixed training with all chunk sizes show the best quality score.
@@ 5.3 Ablation Study
We conduct an ablation study to analyze the mixed training discussed in Sec. 3.2. Specifically, we train models with different mixing schemes as shown in Tab. 3. Chunk size 1 represents the Progressive AR Diffusion [xie2024progressiveautoregressivevideodiffusion] implemented on our model. Recall that our base model has 16 latent frames, so chunk size 16 is the basic T2V. After the models are trained, we generated videos using chunk size $c=1$, since this is the case covered by all models. Among them, we see that the mixture of all chunk sizes achieves the best quality score, although the inference is biased to chunk size 1 as being set in inference. Please note that we use a different inference scheme in Tab. 2, so the numbers of our model are different. The first one with chunk size 1 and no mixing achieves the second-best quality score, indicating an overfitting to this case.
@@ 5.4 Applications
Real-Time Streaming: Our distilled model has a chunk size 2 and 8 sampling steps without CFG. We benchmark its performance on one H100 GPU. It takes 482 ms for one denoising step to generate 2 latent frames and 8 video frames after a $4\times$ TAE. The latency of text encodering and TAE decoding are negligible. Thus, our distilled model can reach real-time performance at 16 FPS. Fig. 7 shows frames from a one-minute long video generated in real-time. With real-time streaming generation, our model can potentially work as a game engine.
@@
Video-to-Video Streaming: Our model can also be applied to real-time video-to-video tasks. The video-to-video process is based on the noise addition and denoising strategy proposed in SDEdit [mengSDEditGuidedImage2022]. First, we add noise to the original input video and then denoise the noisy video with new prompt guidance. We achieve significant content editing while maintaining good temporal consistency. The visual results shown in Fig. 10 highlight the model’s powerful video editing capabilities. In this example, a pig in the input video is modified to a cat in the output video, while the background is preserved.
@@ 7 Acknowledgment
We would like to express our gratitude to Yen-Cheng Liu, Luxin Zhang, and Tao Xu for their valuable discussions and insightful contributions, which greatly enriched our work.
[…]
@@ 10 More Results
To evaluate the capabilities in generating higher-quality videos, we further fine-tuned the 30B model from Movie Gen [moviegen]. While the 30B model is not available for real-time generation at the present, it effectively supports the creation of exceptionally high-quality videos, as illustrated in Fig. 15. Our model consistently produces extended videos characterized by impressive visual quality, content coherence, and diverse scenes aligned accurately with provided text prompts.
@@ 11 Limitations
As described in Sec. 8, the effective context length of StreamDiT fundamentally depends on the base T2V model, and thus lacks long-term memory. When content falls outside the short-term memory window (i.e. context length) of StreamDiT, the associated information is likely to be lost. This can lead to issues such as identity mismatches in a person’s face, or background inconsistencies when the camera makes a full rotation. However, since StreamDiT is orthogonal to additional long-term memory mechanisms, it is possible to address this issue in the future by combining it with long-term memory architectures such as State Space Models like Mamba [gu2023mamba].